# references.bib

### v1

@@ Abstract
In this paper, we consider the resonance problem for the cubic nonlinear Helmholtz equation in the subwavelength regime. We derive a discrete model for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities. Our discrete formulation is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, it characterizes the soliton-like extra eigenmodes that have recently been experimentally observed.
Keywords. Nonlinear subwavelength physics, nonlinear subwavelength resonance, capacitance matrix formalism, cubic nonlinear Helmholtz equation, soliton-like eigenmode.
@@ 1 Introduction and problem formulation
In nanophotonics and nanophononics, the ideal artificial material shall, at subwavelength scales, confine and guide waves, and be able to modify the phase and enhance the amplitude of incoming waves. For doing so, a class of nonlinear artificial materials consisting of nonlinear subwavelength high-contrast resonator building blocks has been introduced. This class of materials has revived the exciting prospect of realizing the long-standing goal of robust transport of waves at subwavelength scales and their amplification in only one direction, opening avenues to channel energy and information transport in quantum computing; see, e.g., [soliton3, rev-acoustics, soliton1, active, soliton2, nonreciprocal].
@@
Although very significant experimental and numerical modeling advances have been achieved in the field of nonlinear subwavelength physics during recent decades, very little is known from a mathematical point of view. As far as we know, properties of nonlinear subwavelength resonances have not yet been analyzed in the mathematical literature. In contrast to the linear response regime studied in depth in [cbms, review], there is a clear lack of deep understanding of the theory of nonlinear subwavelength resonances.
@@
In this work, we lie the basis for nonlinear subwavelength physics. The focus is to develop a theory of nonlinear subwavelength resonances. This is accomplished by using ideas from [feppon2024subwavelength] to the setting of nonlinear systems. It is worth emphasizing that the nonlinearity renders the nonlinear problem much more difficult than the linear one. The total number of subwavelength resonances may be larger than the total number of subwavelength resonators and depends on the strength of the nonlinearities as opposed to the linear systems where the total number of subwavelength resonances is exactly given by the number of resonators. Some soliton-like extra-eigenmodes may exist due to strong enough nonlinearities. These eigenmodes are of profound significance to nonlinear subwavelength physics. They have been experimentally observed in systems of high-contrast resonators where the nonlinearity is of Kerr type [soliton1].
@@
Our main contribution in this work is to provide for the first time a discrete formulation for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities, which is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, our formulation here characterizes the soliton-like extra eigenmodes. Our main tool for building such a discrete framework is a generalization of the Dirichlet-to-Neumann approach for subwavelength resonances from [feppon2024subwavelength] to the nonlinear setting combined with asymptotic analysis of solutions to nonlinear resonance problems for the Helmholtz equation (with respect to the contrast in the material parameters of the resonators).
@@
Concretely, let $\operatorname{{\mathcal{D}}}\subset\operatorname{{\mathbb{R}^{3}}}3$ be a smooth, bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Let $\kappa_{0}\in\mathbb{R}$, $\rho_{0}\in\mathbb{R}$ and let $\kappa_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$, $\rho_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$ be piecewise constant functions. We consider the following system of equations:
@@
$g(u)=b\left|\frac{\partial u}{\partial t}\right|^{2}\frac{\partial u}{\partial t}\quad\text{and}\quad g(u)=b\left(\frac{\partial u}{\partial t}\right)^{2}\overline{u},$
with $b\in\mathbb{C}$. For given operator $g$, we say that $\omega\in\mathbb{C}$ is a resonance of the system (1), if there exists a non-zero $u:\operatorname{{\mathbb{R}^{3}}}3\rightarrow\mathbb{C}$ such that
@@
is a solution to (1). In the linear case, when $g(u)=0$, resonances are precisely those $\omega\in\mathbb{C}$ such that the scattering problem
@@
${c_{0}}:=\sqrt{\frac{\kappa_{0}}{\rho_{0}}},\quad{c_{r}}:=\sqrt{\frac{\kappa_{r}}{\rho_{r}}},\quad\delta(x):=\frac{\rho_{r}(x)}{\rho_{0}}.$
@@
Linear subwavelength resonances have been thoroughly studied since the original paper [ihp]; see, for instance, [cbms, review]. It has been found that the leading-order behavior of subwavelength resonances $\omega(\delta)$ as $\delta\rightarrow 0$ are characterized by the so-called generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$. In fact, one has the following theorem.
Theorem 1.2 (Fundamental theorem of subwavelength physics [review]).
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ be a smooth, bounded domain with well separated connected components $B_{1},\ldots,B_{N}$ which are supposed to have connected boundaries. Then there exists a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ depending on $\operatorname{{\mathcal{D}}}$ and $c_{r}$ and there exists an $\epsilon>0$, such that the following correspondence holds
@@
This theorem has been proved using Gohberg-Sigal theory and layer potential techniques. It lays the basis of many results in subwavelength physics. It has been reproduced using the different approach of Dirichlet-to-Neumann operator by Feppon and Ammari [feppon2024subwavelength] for connected $\operatorname{{\mathcal{D}}}=B_{1}$. In this article, we will enlarge the approach of Dirichlet-to-Neumann operators to disconnected $\operatorname{{\mathcal{D}}}=B_{1}\cup\ldots\cup B_{N}$, reproducing Theorem 1.2, in Section 3. Using similar techniques, we will then develop a theory of nonlinear subwavelength resonances in Section 4.
Let $H^{1}(\operatorname{{\mathcal{D}}})$ denote the usual Sobolev space of square-integrable functions whose weak derivative is square integrable in $\operatorname{{\mathcal{D}}}$. We also denote by $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ the set of traces on $\partial\operatorname{{\mathcal{D}}}$ of functions in $H^{1}(\operatorname{{\mathcal{D}}})$ and let $H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ be the dual of $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$.
@@
Then $\omega$ is a resonance whenever this system is degenerate and does not have a unique solution $u|_{\operatorname{{\mathcal{D}}}}$. Or equivalently $\omega$ is a resonance whenever this system admits a non-zero solution $u|_{\operatorname{{\mathcal{D}}}}$ for $u_{in}\equiv 0$. It was observed in [feppon2024subwavelength] that one can facilitate the analysis and transform this indirect problem of finding characteristic points $\omega$ of the above system into a direct problem, in the case of a connected domain $\operatorname{{\mathcal{D}}}$. By adding $\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the bilinear form on the left and $\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the linear functional on the right one obtains
@@
for $u_{in}\equiv 0$. In the case of sufficiently small $\delta\in\mathbb{R}_{>0}$ and sufficiently small $\omega\in\mathbb{C}$, the bilinear form on the left remains coercive as a small perturbation of the continuous coercive bilinear form [feppon2024subwavelength]
$\begin{split}a_{\omega,\delta}:~{}(u,v)\in H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\longmapsto\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}.\end{split}$
Then the resonances of (2) are characterized as follows [feppon2024subwavelength]: $\omega$ is a resonance if and only if the unique solution $u$ to the variational problem (3) satisfies
@@
allows to solve for $\omega(\delta)$ and determining the subwavelength resonances in the setting of a connected domain $\operatorname{{\mathcal{D}}}$, see [feppon2024subwavelength, Equation (3.21)].
@@
subject to an outgoing radiation condition, where we consider the Kerr-like nonlinearity $\beta\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ inside the resonator domain $\operatorname{{\mathcal{D}}}$. As in the linear case, we will say that $\omega\in\mathbb{C}$ is a resonance of the system (5) whenever it admits a solution of the form $\exp(i\omega t)u(x)$, where $u:\operatorname{{\mathbb{R}^{3}}}3\rightarrow\mathbb{C}$ is non-zero. This allows to consider the following nonlinear Helmholtz resonance problem:
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\rho_{0}}\Delta u+\frac{\omega^{2}}{\kappa_{0}}u&\displaystyle=0&\displaystyle\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{r}(x)}\Delta u+\frac{\omega^{2}}{\kappa_{r}(x)}u&\displaystyle=i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\displaystyle\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\displaystyle\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle u|_{+}&\displaystyle=\left.u\right|_{-}&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (6)
@@
The approach to subwavelength resonances in the nonlinear setting is identical to the approach in the linear setting in Section 3, with the exception that one has to consider a nonlinear form $(u,v)\mapsto\tilde{a}_{\omega,\delta}(u,v)$ for the variational formulation of (6), see Subsection 4.1 and that one has to keep track of the amplitudes of the solutions to the variational problem, see Theorem 4.3. The subwavelength resonances are then determined in the same way as for the linear setting, by expanding in $\omega$ and $\delta$ and determining $\omega(\delta)$ such that the characterization of a subwavelength frequency is satisfied, see Subsection 4.2.
This article is organized as follows. In Section 2 we set the notation and the technical prerequisites which will be of use in the later sections. Section 3 is dedicated to the characterization of subwavelength resonances of linear Helmhotz systems with disconnected resonator domain $\operatorname{{\mathcal{D}}}$. It is divided into Subsection 3.1 where the partial differential equation (PDE) system is formulated variationally and the resonances are variationally characterized and into Subsection 3.2 where the asymptotics of subwavelength resonances are computed. Section 4 then builds on top of these techniques and follows a similar structure. First in Subsection 4.1 the system of nonlinear PDE’s is formulated variationally and resonances are characterized. Then in Subsection 4.2 the asymptotics of nonlinear subwavelength resonances are derived. In Section 5 simulations of the leading-order asymptotics of the subwavelength resonances of a nonlinear system of dimers $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ are presented, demonstrating the existence of more subwavelength resonances and more subwavelength solutions than in the linear case. This article is concluded with Section 6 discussing final conclusions and an outlook for further research.

### v3

# Nonlinear subwavelength resonances in three dimensions

@@ Abstract
In this paper, we consider the resonance problem for the cubic nonlinear Helmholtz equation in the subwavelength regime. We derive a discrete model for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities. Our discrete formulation is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, it characterizes the experimentally observed extra eigenmodes that are induced by the nonlinearities.
Keywords. Nonlinear subwavelength physics, nonlinear subwavelength resonance, capacitance matrix formalism, cubic nonlinear Helmholtz equation, nonlinearity-induced eigenmode.
@@ 1 Introduction and problem formulation
In nanophotonics and nanophononics, the ideal artificial material shall, at subwavelength scales, confine and guide waves, and be able to modify the phase and enhance the amplitude of incoming waves. In order to do so, a class of nonlinear artificial materials consisting of nonlinear subwavelength high-contrast resonator building blocks has been introduced. This class of materials has revived the exciting prospect of realizing the long-standing goal of robust transport of waves at subwavelength scales and their amplification in only one direction, opening avenues to channel energy and information transport in quantum computing; see, e.g., [10, 16, 19, 22, 24, 25].
@@
Although very significant advances in experimental and numerical modeling have been achieved in the field of nonlinear subwavelength physics during recent decades, very little is known from a mathematical point of view. To our knowledge, the properties of nonlinear subwavelength resonances have not yet been analyzed in the mathematical literature. In contrast to the linear response regime studied in depth in [7, 8], there is a clear lack of deep understanding of the theory of nonlinear subwavelength resonances.
@@
In this work, we lie the basis for nonlinear subwavelength physics. The focus is to develop a theory of nonlinear subwavelength resonances. This is accomplished by using ideas from [12] in the context of nonlinear systems. It is worth emphasizing that the nonlinearity renders the nonlinear problem much more difficult than the linear one. The total number of subwavelength resonances may be larger than the total number of subwavelength resonators and depends on the strength of the nonlinearities as opposed to the linear systems where the total number of subwavelength resonances is exactly given by the number of resonators. Some extra-eigenmodes may exist due to strong enough nonlinearities. These eigenmodes are of profound significance to nonlinear subwavelength physics. They have been experimentally observed in systems of high-contrast resonators where the nonlinearity is of Kerr type [19].
@@
Our main contribution in this work is to provide for the first time a discrete formulation for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities, which is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, our formulation here characterizes the extra eigenmodes induced by the nonlinearities. Our main tool for building such a discrete framework is a generalization of the Dirichlet-to-Neumann approach for subwavelength resonances from [12] to the nonlinear setting combined with asymptotic analysis of solutions to nonlinear resonance problems for the Helmholtz equation (with respect to the contrast in the material parameters of the resonators).
@@
Concretely, let $\operatorname{{\mathcal{D}}}\subset\operatorname{{\mathbb{R}^{3}}}3$ be a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Let $\kappa_{0}\in\mathbb{R}$, $\rho_{0}\in\mathbb{R}$ and let $\kappa_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$, $\rho_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$ be piecewise constant functions. We consider the following system of equations:
@@
$g(u)={\frac{\beta}{\kappa_{r}}}\left|\frac{\partial u}{\partial t}\right|^{2}\frac{\partial u}{\partial t}\quad\text{and}\quad g(u)=\frac{\beta}{\kappa_{r}}\left(\frac{\partial u}{\partial t}\right)^{2}\overline{u},$
with $\beta\in\mathbb{C}$. For given operator $g$, we say that $\omega\in\mathbb{C}$ is a resonance of the system (1), if there exists a non-zero $u:\operatorname{{\mathbb{R}^{3}}}3\rightarrow\mathbb{C}$ such that
@@
is a solution to (1).
In the linear case, when $g(u)=0$, resonances are precisely those $\omega\in\mathbb{C}$ such that the scattering problem
@@
${c_{0}}:=\sqrt{\frac{\kappa_{0}}{\rho_{0}}},\quad{c_{r}}(x):=\sqrt{\frac{\kappa_{r}(x)}{\rho_{r}(x)}},\quad\delta(x):=\frac{\rho_{r}(x)}{\rho_{0}}.$
@@
Linear subwavelength resonances have been thoroughly studied since the original paper [4]; see, for instance, [7, 8]. The leading-order behavior of subwavelength resonances $\omega(\delta)$ as $\delta\rightarrow 0$ arecharacterized by the so-called generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$. In fact, one has the following theorem.
Theorem 1.2 (Fundamental theorem of subwavelength physics [8]).
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ be a smooth bounded domain with well-separated connected components $B_{1},\ldots,B_{N}$ which are supposed to have connected boundaries. Then there exists a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ depending on $\operatorname{{\mathcal{D}}}$ and $c_{r}$ and there exists an $\epsilon>0$, such that the following correspondence holds
@@
This theorem has been proved using the Gohberg-Sigal theory and layer potential techniques. It sets the basis of many results in subwavelength physics. It has been reproduced using the different approach of Dirichlet-to-Neumann operator by Feppon and Ammari [12] for connected $\operatorname{{\mathcal{D}}}=B_{1}$. In this article, we will enlarge the approach of Dirichlet-to-Neumann operators to disconnected $\operatorname{{\mathcal{D}}}=B_{1}\cup\ldots\cup B_{N}$, reproducing Theorem 1.2, in Section 3. Using similar techniques, we will then develop a theory of nonlinear subwavelength resonances in Section 4.
Let $H^{1}(\operatorname{{\mathcal{D}}})$ denote the usual Sobolev space of square-integrable functions whose weak derivative is square-integrable in $\operatorname{{\mathcal{D}}}$. We also denote by $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ the set of traces on $\partial\operatorname{{\mathcal{D}}}$ of the functions in $H^{1}(\operatorname{{\mathcal{D}}})$ and let $H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ be the dual of $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$.
@@
Then $\omega$ is a resonance whenever this system is degenerate and does not have a unique solution $u|_{\operatorname{{\mathcal{D}}}}$. Or equivalently $\omega$ is a resonance whenever this system admits a non-zero solution $u|_{\operatorname{{\mathcal{D}}}}$ for $u_{in}\equiv 0$. It was observed in [12] that one can facilitate the analysis and transform this indirect problem of finding characteristic points $\omega$ of the above system into a direct problem, in the case of a connected domain $\operatorname{{\mathcal{D}}}$. By adding $\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the bilinear form on the left and $\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the linear functional on the right one obtains
@@
for $u_{in}\equiv 0$. In the case of sufficiently small $\delta\in\mathbb{R}_{>0}$ and sufficiently small $\omega\in\mathbb{C}$, the bilinear form on the left remains coercive as a small perturbation of the continuous coercive bilinear form
$\begin{split}a_{\omega,\delta}:~{}(u,v)\in H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\longmapsto\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}\end{split}$
does not change its coercivity [12]. Then the resonances of (2) are characterized as follows [12]: $\omega$ is a resonance if and only if the unique solution $u$ to the variational problem (3) satisfies
@@
allows to solve for $\omega(\delta)$ and to determine asymptotically the subwavelength resonances in the setting of a connected domain $\operatorname{{\mathcal{D}}}$, see [12, Equation (3.21)].
@@
subject to an outgoing radiation condition, where we consider the Kerr-like nonlinearity $\beta\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ inside the resonator domain $\operatorname{{\mathcal{D}}}$. As in the linear case, we will say that $\omega\in\mathbb{C}$ is a resonance of the system (5) whenever it admits a non-zero solution of the form $\exp(i\omega t)u(x)$. This allows to consider the following nonlinear Helmholtz resonance problem:
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\rho_{0}}\Delta u+\frac{\omega^{2}}{\kappa_{0}}u&\displaystyle=0&\displaystyle\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{r}(x)}\Delta u+\frac{\omega^{2}}{\kappa_{r}(x)}u&\displaystyle={-}i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\displaystyle\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\displaystyle\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle u|_{+}&\displaystyle=\left.u\right|_{-}&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (6)
@@
The approach to subwavelength resonances in the nonlinear setting is identical to the approach in the linear setting in Section 3, with the exception that one has to consider a nonlinear form $(u,v)\mapsto\tilde{a}_{\omega,\delta}(u,v)$ for the variational formulation of (6), see Subsection 4.1 and that one has to keep track of the amplitudes of the solutions to the variational problem, see Theorem 4.3. The subwavelength resonances are then determined in the same way as for the linear setting, by expanding in $\omega$ and $\delta$ and determining $\omega(\delta)$ such that the characterization of a subwavelength frequency is satisfied; see Subsection 4.2.
This article is organized as follows. In Section 2 we set the notation and technical prerequisites which will be useful in the later sections. Section 3 is dedicated to the characterization of subwavelength resonances of linear Helmholtz systems with disconnected resonator domain $\operatorname{{\mathcal{D}}}$. It is divided into Subsection 3.1 where the partial differential equation (PDE) system is formulated variationally and the resonances are variationally characterized and into Subsection 3.2 where the asymptotics of subwavelength resonances are computed. Section 4 then builds on top of these techniques and follows a similar structure. First in Subsection 4.1 the system of nonlinear PDE’s is formulated variationally and resonances are characterized. Then, in Subsection 4.2, the asymptotics of nonlinear subwavelength resonances are derived. In Section 5 simulations of the leading-order asymptotics of the subwavelength resonances of a nonlinear system of dimers $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ are presented, demonstrating the existence of more subwavelength resonances and more subwavelength solutions than in the linear case. This article is concluded with Section 6 discussing the final conclusions and an outlook for further research.